### v1

@@ Abstract
As autonomous systems are increasingly deployed in open and uncertain settings, there is a growing need for trustworthy world models that can reliably predict future high-dimensional observations. The learned latent representations in world models lack direct mapping to meaningful physical quantities and dynamics, limiting their utility and interpretability in downstream planning, control, and safety verification. In this paper, we argue for a fundamental shift from physically informed to physically interpretable world models — and crystallize four principles that leverage symbolic knowledge to achieve these ends: (1) structuring latent spaces according to the physical intent of variables, (2) learning aligned invariant and equivariant representations of the physical world, (3) adapting training to the varied granularity of supervision signals, and (4) partitioning generative outputs to support scalability and verifiability. We experimentally demonstrate the value of each principle on two benchmarks. This paper opens several intriguing research directions to achieve and capitalize on full physical interpretability in world models.
@@ 1 Introduction
Autonomous systems are increasingly deployed in open and uncertain environments [Saidi et al.(2022)Saidi, Ziegenbein, Deshmukh, and Ernst, Topcu et al.(2020)Topcu, Bliss, Cooke, Cummings, Llorens, Shrobe, and Zuck] and use high-dimensional observations like high-resolution images and LiDAR scans to perceive these environments in necessary detail. To achieve high performance, their planning and control are often implemented with deep learning methods like reinforcement learning (RL) [Yang et al.(2022)Yang, Zhang, Luu, Ha, Tan, and Yu, Garg et al.(2019)Garg, Chiang, Sugaya, Faust, and Tapia]. Since RL training is sample-inefficient, it is impractical to perform in the real world — leading the controllers to be trained “in the imagination” of world models [Ha and Schmidhuber(2018b), Wu et al.(2022)Wu, Escontrela, Hafner, Abbeel, and Goldberg].
World models learn to approximate the physical world by predicting future observations based on current observations and actions. Popular neural world models rely on compressing high-dimensional observations into the latent space using an autoencoder. Then these latent values are propagated forward in time based on learned temporal dependencies [Deng et al.(2023)Deng, Park, and Ahn] and decoded into predicted observations. World models can be drastically improved by injecting symbolic physical knowledge into their structure and training process. For example, the approach by [Chen et al.(2022)Chen, Huang, Raghupathi, Chandratreya, Du, and Lipson] automatically discovers physically meaningful variables from raw observations for more stable long-horizon predictions than methods with standard high-dimensional autoencoders. In another instance, [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis] decompose control into neuro-symbolic predicates powered by vision-language models and predefined control primitives, thereby improving generalization.
A major challenge of modern world models is their lack of physical interpretability: usually the latent state cannot be easily mapped to physical quantities (e.g., pose or velocity). This limits their use in classical model-based autonomy and the design of physically-grounded rewards for RL. We also cannot obtain physical guarantees from reachability analysis based on world models [Katz et al.(2022)Katz, Corso, Strong, and Kochenderfer]. The core reason for this uninterpretability is that deep learning thrives on distributed representations, in which each feature is partially encoded in multiple latent variables [Hinton(1986)]. This challenge is further complicated by partial online observability of the physical state and the difficulty of precisely labeling the data (e.g., indicating which state is riskier in a video).
@@
Principle 1: …be structured according to the physical intent of latent variables.
@@
Principle 3: …adapt their training to the varied granularity of supervision signals.
@@ 2 World Models: State of the Art
Foundations of world models. Modern world models have led to state-of-the-art performance in autonomous planning and control while addressing the data-efficiency concerns of standard RL [Deng et al.(2023)Deng, Park, and Ahn, Micheli et al.(2023b)Micheli, Alonso, and Fleuret, Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling]. Modern deep-world models were first proposed by [Ha and Schmidhuber(2018a)]: a variational autoencoder (VAE) generated image representations and a recurrent neural network made latent-space predictions. Later work refined both the encoder-decoder architecture and the surrogate dynamics: PlaNet introduced a recurrent state-space model (RSSM) for prediction [Hafner et al.(2018)Hafner, Lillicrap, Fischer, Villegas, Ha, Lee, and Davidson]; Dreamer propagated gradients back through the imagined trajectories to refine latent prediction [Hafner et al.(2019b)Hafner, Lillicrap, Ba, and Norouzi]; and DreamerV2 extended the RSSM to categorical latent variables [Hafner et al.(2022)Hafner, Lillicrap, Norouzi, and Ba]. More recent research integrates autoregressive transformers with self-attention layers to capture detailed temporal dependencies [Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling], or diffusion models to mitigate compounding errors [Ding et al.(2024)Ding, Zhang, Tian, and Zheng]. World models have also found success in optimizing planning algorithms for autonomous vehicles in realistic environments: DriveDreamer [Wang et al.(2023b)Wang, Zhu, Huang, Chen, Zhu, and Lu] leverages multi-modal inputs to generate realistic video trajectories for policy optimization; DriveWorld [Min et al.(2024)Min, Zhao, Xiao, Zhao, Xu, Zhu, Jin, Li, Guo, Xing, Jing, Nie, and Dai], OccWorld [Zheng et al.(2023)Zheng, Chen, Huang, Zhang, Duan, and Lu], UniWorld [Min et al.(2023)Min, Zhao, Xiao, Nie, and Dai], and RenderWorld [Yan et al.(2024)Yan, Dong, Shao, Lu, Haiyang, Liu, Wang, Wang, Wang, Remondino, and Ma] forecast detailed 3D occupancy to inform motion planning and control.
@@
Towards interpretable world models. Despite the strong performance of world models, their interpretability remains a major challenge in most frameworks. Early efforts toward disentangling latent variables (i.e., reducing their mutual dependency) include -VAEs [Higgins et al.(2016)Higgins, Matthey, Pal, Burgess, Glorot, Botvinick, Mohamed, and Lerchner] and causal VAEs [Yang et al.(2021)Yang, Liu, Chen, Shen, Hao, and Wang]. This disentanglement strategy is also employed in driving prediction frameworks like GNeVA [Lu et al.(2024a)Lu, Zhan, Tomizuka, and Hu] and ISAP [Itkina and Kochenderfer(2022)]. Under the umbrella of world models, G-SWM [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] investigated a principled modeling framework that inherits interpretable object and context latent separation from various spatial attention approaches [Kosiorek et al.(2018)Kosiorek, Kim, Posner, and Teh, Kossen et al.(2020)Kossen, Stelzner, Hussing, Voelcker, and Kersting, Jiang et al.(2020)Jiang, Janghorbani, Melo, and Ahn, Crawford and Pineau(2020)]. Interpretability can also manifest in the world model’s forward dynamics: [Cui et al.(2020)Cui, Nguyen, Chou, Lin, Schneider, Bradley, and Djuric] propose embedding known dynamics into neural pipelines, while more recent work imposes further physical constraints on system identification [Sridhar et al.(2023)Sridhar, Dutta, Weimer, and Lee], motion prediction [Tumu et al.(2023)Tumu, Lindemann, Nghiem, and Mangharam], and ordinary differential equations with learnable parameters [Linial et al.(2021)Linial, Ravid, Eytan, and Shalit, Zhong and Meidani(2023), Mao et al.(2025)Mao, Gu, Sha, Shao, Wang, and Abdelzaher]. Incorporating partial knowledge of physics with weak supervision has recently improved both the state and dynamics interpretability [Mao and Ruchkin(2024)]. A recent Nature article leveraged the biological alignment of latent representations to predict microbiome community interactions and antibiotic resistance [Baig et al.(2023)Baig, Ma, Xu, and You]. Neuro-symbolic world models have also begun to emerge: VisualPredicator [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis] learns a set of abstract states and high-level actions for strong out-of-distribution generalization, whereas WorldCloner [Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl] learns symbolic rules to adapt the dynamics to open world novelty.
@@
Knowledge gap. We observe the lack of world model architectures supporting full physical interpretability, as illustrated in Figure 1 (the underlying literature list can be found in Table 2 in the appendix). Some existing neuro-symbolic architectures scrape the threshold of physically interpretable dynamics, yet lack fluid state representations. On the other hand, powerful multimodal transformer-based architectures preserve the physical context through 3D occupancy, but their prediction mechanisms are black-box. Bridging this gap is key to transitioning from merely physically informed world models to fully physically interpretable ones. Our recent work highlights the need to address these open questions [Lu et al.(2024b)Lu, Afridi, Kang, Ruchkin, and Zheng] and the promise of predictive world models [Mao et al.(2024b)Mao, Sobolewski, and Ruchkin] and their foundation-model variants [Mao et al.(2024a)Mao, Dai, Geng, and Ruchkin].
@@
Benefits of Physical Interpretability. Aligning world models with fundamental physical principles (e.g., kinematics and conservation laws) has been shown to improve their out-of-distribution generalization and robustness [Mao and Ruchkin(2024), Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn, Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis, Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl]. These principles prevent latching onto spurious correlations in training and constrain the models to traverse a physically meaningful manifold when extrapolating observations. Going further, physically interpretable representations would lead us to a qualitatively new level of safety and trustworthiness. It would make world models more transparent and debuggable by cross-checking them with real-world physics. It would also make generative components suitable for closed-loop verification of physical properties. Finally, physical representation would drastically improve RL sample efficiency by shrinking the search space to physically feasible solutions.
@@ 3 Physical Interpretability Principles for World Models
3.1 Principle 1: Structuring the Latent Space with Prior Knowledge
Although many facets of human cognition remain a mystery, extensive research indicates that distinct regions of the brain are dedicated to unique functions [Luria(1970)]. For example, early studies of patients with brain injuries revealed the existence of functional systems partially dedicated to meaningful tasks such as spatial awareness, voluntary movement, and sensory processing [Luria(1970)]. While modeling the minutia of human cognition is beyond the scope of this principle, this idea of functional organization inspires how world model architectures are increasingly structured.
@@ 3.1 Principle 1: Structuring the Latent Space with Prior Knowledge
Recent work in the context of Branch 1 aligned latent representations with physical properties [Mao and Ruchkin(2024)]. Although modeling agent interactions is a well-explored topic, early work by [Kipf et al.(2018)Kipf, Fetaya, Wang, Welling, and Zemel] revealed how physical interactions between agents can be learned through graph neural networks (GNN) in an unsupervised manner. Later advances in learning physical interactions within the context of world models by [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] proposed a method that constructs a separate latent representation using a GNN to capture agent occlusion and interaction relationships. Finally, Branch 3 follows the typical strategy for creating uninterpretable world models and is considered a useful layer to the structured latent space [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn].
Latent space structuring has become increasingly prevalent in improving the performance of planning and control. To this end, [Lu et al.(2024a)Lu, Zhan, Tomizuka, and Hu] developed a goal-based neural variational agent (GNeVA) that uses separate polyline embeddings for the agent and the map to achieve interpretable generative motion prediction. Similarly, [Itkina and Kochenderfer(2022)] designed an interpretable car trajectory prediction that integrates three distinct workflow branches: agent states, high-definition maps, and social context.
@@ 3.2 Principle 2: Exploiting Invariances and Equivariances in Input and Latent Spaces
Deep neural networks have achieved impressive performance due in part to their ability to learn rich distributed representations from training data. Rather than simply memorizing examples, these models construct hierarchical feature embeddings that capture patterns in the data to generalize to i.i.d. samples [Hinton(1986)]. Nevertheless, training a model to internalize and imagine the world in a human-like manner from scratch is still a nontrivial challenge at best ([Ha and Schmidhuber(2018a)]). Encoding high-dimensional observations (e.g., images) through commonplace embedding methods (e.g., through autoencoders or encoder-only transformers) leaves the latent representation generally uninterpretable and task-agnostic. This raises concerns about whether the latent space is distorted by spurious correlations or if it effectively encodes the details necessary for discriminating between features that should remain functionally disentangled.
Invariance and equivariance relations are one way to address uninterpretability in representation learning models. These terms are often used to classify representations based on their response to observation space transformations. If the representation of shifts in an expected manner due to a transformation , then the representation model is said to be equivariant to that transformation. Likewise, if the representation does not shift under the transformation, then the model is said to be invariant to the transformation.
@@
To this end, [Zhang et al.(2020)Zhang, Lyle, Sodhani, Filos, Kwiatkowska, Pineau, Gal, and Precup] utilize bisimulation metrics to learn latent representations of roadway obstacles that are invariant to obstacle type, size, and brightness. The dual of avoiding explicitly irrelevant details is explicitly learning relevant details. [Pol et al.(2020)Pol, Kipf, Oliehoek, and Welling] introduce contrastive loss to enforce action-equivariance on learned representations. However, we still lack ways to effectively incorporate these expert conceptual priors into world models. [Park et al.(2022)Park, Biza, Zhao, Meent, and Walters] innovate a symmetric embedding network for world models that learns simple latent space transformations from complex observation space transformations.
@@
We classify representation models along two dimensions: (1) the nature of their transformation response (invariance versus equivariance) and (2) their degree of human alignment (aligned versus misaligned). A representation that is aligned-invariant remains unchanged when an observation undergoes a meaning-preserving transformation, while an aligned-equivariant representation transforms predictably when the observation’s meaning is altered. In contrast, a representation is misaligned-invariant if it does not change under meaningful effects made to the observation (suggesting underfitting), and it is misaligned-equivariant if it changes in response to an observation transformation that should not affect the underlying meaning (suggesting a domain shift). Our training objective is to achieve invariance and equivariance alignment by ensuring that the post-transformation representations accurately reflect our human interpretation of the change.
Definition 3.1 (Equivariant Representation).
@@
Let be a function that maps each observation to a latent representation . Let be a finite set of transformation pairs such that and . We say that is equivariant to if . Invariance is a special case of equivariance where .
@@
In practice, deciding which features the representation should be invariant or equivariant to solely depends on the task at hand and the discretion of the expert engineer. In autonomous driving, consider a recovery component that overrides the default controller during inclement weather. This component’s representations of real-time images might change when the sky is cloudy. On the other hand, the world model of the default controller might be invariant to the features of the sky.
3.3 Principle 3: Multi-Level Supervision for Latent Representations
World models must bridge the gap between rich observations and physical meaning. This requires adapting the training strategy based on the type and quality of available supervision signals [Lee et al.(2013), Chen et al.(2020)Chen, Kornblith, Norouzi, and Hinton]. Instead of treating all supervision signals in the same way, effective world models carefully exploit supervision signals — whether they are precise, coarse, or missing — by selecting the most fitting training method to align with the underlying physical system.
Multi-level supervision tailors the loss functions and training process to the granularity of supervision signals. For instance, physical state labels allow for the direct alignment of latent representations with real-world quantities using supervised loss. When such labels are unavailable, temporal consistency and smoothness of trajectories can serve as implicit regularization techniques to constrain learned representations. Finally, self-supervision can leverage data-driven structures to discover meaningful latent representations in entirely unsupervised settings.
Supervised Learning: Direct supervision connects latent values with physical states. In many cases, supervision signals are introduced directly into the embeddings to capture key features from labeled data [Zhuang et al.(2015)Zhuang, Cheng, Luo, Pan, and He]. For example, in low-dimensional systems with position and velosity states , additional latent dimensions () can improve reconstruction quality and stability [Chen et al.(2016)Chen, Duan, Houthooft, Schulman, Sutskever, and Abbeel, Alemi et al.(2018)Alemi, Poole, Fischer, Dillon, Saurous, and Murphy, Rezende et al.(2014)Rezende, Mohamed, and Wierstra].

### v2

@@ Abstract
As autonomous systems are increasingly deployed in open and uncertain settings, there is a growing need for trustworthy world models that can reliably predict future high-dimensional observations. The learned latent representations in world models lack direct mapping to meaningful physical quantities and dynamics, limiting their utility and interpretability in downstream planning, control, and safety verification. In this paper, we argue for a fundamental shift from physically informed to physically interpretable world models — and crystallize four principles that leverage symbolic knowledge to achieve these ends: (1) functionally organizing the latent space according to the physical intent, (2) learning aligned invariant and equivariant representations of the physical world, (3) integrating multiple forms and strengths of supervision into a unified training process, and (4) partitioning generative outputs to support scalability and verifiability. We experimentally demonstrate the value of each principle on two benchmarks. This paper opens several intriguing research directions to achieve and capitalize on full physical interpretability in world models.
@@ 1 Introduction
Autonomous systems are increasingly deployed in open and uncertain environments [Saidi et al.(2022)Saidi, Ziegenbein, Deshmukh, and Ernst, Topcu et al.(2020)Topcu, Bliss, Cooke, Cummings, Llorens, Shrobe, and Zuck] and use high-dimensional observations to perceive these environments in necessary detail. To achieve high performance, planning and control are often implemented with deep learning methods like reinforcement learning (RL) [Yang et al.(2022)Yang, Zhang, Luu, Ha, Tan, and Yu, Garg et al.(2019)Garg, Chiang, Sugaya, Faust, and Tapia]. Since RL training is sample-inefficient, it is impractical to perform in the real world — leading to controllers trained “in the imagination” of world models [Ha and Schmidhuber(2018), Wu et al.(2022)Wu, Escontrela, Hafner, Abbeel, and Goldberg].
World models learn to approximate the physical world by predicting future observations based on current observations and actions. Popular neural world models compress observations into the latent space using an autoencoder, propagate these latent values forward in time based on learned temporal dependencies [Deng et al.(2023)Deng, Park, and Ahn], and decode them into predicted observations. World models can be improved by injecting symbolic physical knowledge into their structure and training process. For example, [Chen et al.(2022)Chen, Huang, Raghupathi, Chandratreya, Du, and Lipson] automatically extracted physically meaningful variables from raw observations, yielding more stable long-horizon predictions than standard autoencoders. [Brunton et al.(2016)Brunton, Proctor, and Kutz] similarly used sparse regression to recover governing equations of nonlinear dynamics from noisy data. Controllers also generalize better when expressed as neuro-symbolic predicates that combine vision-language models with predefined control primitives [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis].
A major challenge of modern world models is their lack of physical interpretability. We define it as the degree to which a model’s learned latent space corresponds meaningfully to the underlying physics: (a) how well latent embeddings map to physical variables, and (b) how closely latent dynamics emulate physical processes. Without sufficient physical interpretability, a world model offers limited utility in classical model-based autonomy and the design of physically grounded rewards for RL. We also cannot obtain physical guarantees from reachability analysis based on world models [Katz et al.(2022)Katz, Corso, Strong, and Kochenderfer]. The core reason for this uninterpretability is that deep learning thrives on distributed representations, in which each feature is partially encoded in multiple latent variables [Hinton(1986)]. This challenge is further complicated by partial online observability of the physical state and the difficulty of precisely labeling the data (e.g., indicating which state is riskier in a video).
@@
Principle 1: …have a functionally organized latent space.
@@
Principle 3: …integrate multiple forms and strengths of supervision into training.
@@ 2 World Models: State of the Art
Foundations of world models. Modern world models have led to state-of-the-art performance in autonomous planning and control while addressing the data-efficiency concerns of standard RL [Deng et al.(2023)Deng, Park, and Ahn, Micheli et al.(2023)Micheli, Alonso, and Fleuret, Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling]. Early world models combined a variational autoencoder (VAE) with a recurrent neural network to predict latent dynamics [Ha and Schmidhuber(2018)]. Later work refined both the encoder-decoder architecture and the surrogate dynamics: PlaNet introduced a recurrent state-space model (RSSM) for prediction [Hafner et al.(2019)Hafner, Lillicrap, Fischer, Villegas, Ha, Lee, and Davidson]; Dreamer backpropagated gradients through imagined trajectories to improve latent prediction [Hafner et al.(2020)Hafner, Lillicrap, Ba, and Norouzi]; and DreamerV2 extended the RSSM to categorical latent variables [Hafner et al.(2022)Hafner, Lillicrap, Norouzi, and Ba]. More recent research combines autoregressive transformers with self-attention layers to capture detailed temporal dependencies [Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling], or diffusion models to mitigate compounding errors [Ding et al.(2024)Ding, Zhang, Tian, and Zheng]. World models have been used to optimize planning algorithms for autonomous vehicles in realistic environments: DriveDreamer [Wang et al.(2023b)Wang, Zhu, Huang, Chen, Zhu, and Lu] generates realistic video trajectories from multi-modal inputs for policy optimization; DriveWorld [Min et al.(2024)Min, Zhao, Xiao, Zhao, Xu, Zhu, Jin, Li, Guo, Xing, Jing, Nie, and Dai], OccWorld [Zheng et al.(2023)Zheng, Chen, Huang, Zhang, Duan, and Lu], UniWorld [Min et al.(2023)Min, Zhao, Xiao, Nie, and Dai], and RenderWorld [Yan et al.(2024)Yan, Dong, Shao, Lu, Haiyang, Liu, Wang, Wang, Wang, Remondino, and Ma] forecast detailed 3D occupancy for motion planning.
@@
Towards interpretable world models. Despite the strong performance of world models, their interpretability remains a challenge in most frameworks. Relevant efforts toward disentangling latent variables (i.e., reducing their mutual dependency) include -VAEs [Higgins et al.(2016)Higgins, Matthey, Pal, Burgess, Glorot, Botvinick, Mohamed, and Lerchner] and causal VAEs [Yang et al.(2021)Yang, Liu, Chen, Shen, Hao, and Wang]. This disentanglement strategy is also employed in driving prediction frameworks like GNeVA [Lu et al.(2024a)Lu, Zhan, Tomizuka, and Hu] and ISAP [Itkina and Kochenderfer(2022)]. Under the umbrella of world models, G-SWM [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] investigated a principled modeling framework that inherits interpretable object and context latent separation from various spatial attention approaches [Kosiorek et al.(2018)Kosiorek, Kim, Posner, and Teh, Kossen et al.(2020)Kossen, Stelzner, Hussing, Voelcker, and Kersting, Jiang et al.(2020)Jiang, Janghorbani, Melo, and Ahn, Crawford and Pineau(2020)]. [Fremont et al.(2019)Fremont, Dreossi, Ghosh, Yue, Sangiovanni-Vincentelli, and Seshia] proposed SCENIC — a probabilistic program for generating realistic scenes with physical constraints. More recent methods impose physical constraints for system identification [Sridhar et al.(2023)Sridhar, Dutta, Weimer, and Lee], motion prediction [Tumu et al.(2023)Tumu, Lindemann, Nghiem, and Mangharam], and learnable ODE modeling [Linial et al.(2021)Linial, Ravid, Eytan, and Shalit, Zhong and Meidani(2023), Mao et al.(2025)Mao, Gu, Sha, Shao, Wang, and Abdelzaher]. Incorporating partial knowledge of physics with weak supervision has also improved both the state and dynamics interpretability [Mao and Ruchkin(2024)]. A recent Nature article leveraged the biological alignment of latent representations to predict microbiome community interactions and antibiotic resistance [Baig et al.(2023)Baig, Ma, Xu, and You]. Neuro-symbolic world models have also begun to emerge: VisualPredicator [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis] learns a set of abstract states and high-level actions for strong out-of-distribution generalization, whereas WorldCloner [Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl] learns symbolic rules to adapt the dynamics to open world novelty. Relevant neuro-symbolic research includes PhysORD [Zhao et al.(2024)Zhao, Li, Du, Fu, and Wang], which embeds physical laws into neural models, and work by [Miao et al.(2025)Miao, Fainekos, Hoxha, Okamoto, Prokhorov, and Mitra] transforming dashcam footage from a driving environment into a SCENIC script through a vision-language model.
@@
Knowledge gap. We observe the lack of world models with full physical interpretability, as per Figure 1 (the underlying literature is listed in Table 2 in the Appendix). Some existing neuro-symbolic architectures scrape the threshold of physically interpretable dynamics, yet lack fluid state representations. On the other hand, multimodal transformer-based architectures preserve the physical context through 3D occupancy but predict with black-box mechanisms. Bridging this gap is key to transitioning from merely physically informed world models to fully physically interpretable ones. Our recent work highlights the need to address these open questions [Lu et al.(2024b)Lu, Afridi, Kang, Ruchkin, and Zheng] with predictive world models [Mao et al.(2024b)Mao, Sobolewski, and Ruchkin] and their foundation-model variants [Mao et al.(2024a)Mao, Dai, Geng, and Ruchkin].
@@
Benefits of Physical Interpretability. Aligning world models with fundamental physical principles (e.g., kinematics and conservation laws) has been shown to improve their out-of-distribution generalization and robustness [Mao and Ruchkin(2024), Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn, Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis, Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl, Greydanus et al.(2019)Greydanus, Dzamba, and Yosinski]. These principles prevent latching onto spurious correlations in training and constrain the models to traverse a physically meaningful manifold when extrapolating observations. Going further, physically interpretable representations would lead us to a qualitatively new level of safety and trustworthiness. It would make world models more transparent and debuggable by cross-checking them with real-world physics. It would also make generative components suitable for closed-loop verification of physical properties. Finally, physical representation would improve RL sample efficiency by shrinking the search space to physically feasible solutions.
@@ 3 Physical Interpretability Principles for World Models
This section puts forward four guiding principles for building physically interpretable world models. We begin by formally defining a world model:
Definition 3.1 (World Model).
A world model is a function that maps an observation to , where is the discrete time index, maps the observation to a latent embedding , is the latent dynamics propagating to , and maps embedding to observation .
A world model learns its latent space by minimizing the gap between predicted and actual observations. However, due to being a black box, its latent space often lacks direct physical interpretability. To address this, we introduce a general concept of a physically interpretable world model:
Definition 3.2 (Physically Interpretable World Model).
A world model is physically interpretable if: (i) there exists a latent-to-physical mapping , where is the degrees of freedom of the physical environment, such that , where is the minimal set of state variables required to fully describe the environment’s true dynamics ; and (ii) it holds that: for all .
3.1 Principle 1: Functionally Organizing the Latent Space with Prior Knowledge
@@
Recent work for the first branch aligned latent representations with physical properties [Mao and Ruchkin(2024)]. For the second branch, earlier studies demonstrated that physical interactions between agents can be learned without supervision through graph neural networks (GNNs) [Kipf et al.(2018)Kipf, Fetaya, Wang, Welling, and Zemel]. In the context of world models, [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] constructed a separate latent representation using a GNN to capture agent occlusions and interactions. Physics-informed neural networks [Raissi et al.(2019)Raissi, Perdikaris, and Karniadakis, Saemundsson et al.(2020)Saemundsson, Terenin, Hofmann, and Deisenroth] can also improve the physical interpretability of the world model’s dynamics. For instance, Hamiltonian neural networks [Greydanus et al.(2019)Greydanus, Dzamba, and Yosinski] learn and adhere to physical conservation laws, leading to impressive generalization. The third branch follows the typical strategy for creating uninterpretable world models and is considered a useful layer to the structured latent space [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn]. Latent space structuring has become increasingly prevalent to improve the performance of planning and control. For instance, a goal-based neural variational agent (GNeVA) uses separate polyline embeddings for the agent and the map, enabling interpretable generative motion prediction [Lu et al.(2024a)Lu, Zhan, Tomizuka, and Hu]. Similarly, an interpretable car trajectory prediction framework was proposed, integrating three distinct workflow branches: agent states, high-definition maps, and social context [Itkina and Kochenderfer(2022)].
@@ 3.2 Principle 2: Exploiting Invariances and Equivariances in Input and Latent Spaces
Neural networks’ impressive performance is due in part to their ability to learn rich distributed representations from training data. Rather than memorizing examples, these models construct hierarchical feature embeddings that capture data patterns and generalize to i.i.d. samples [Hinton(1986)]. Nevertheless, training a model to internalize and imagine the world in a human-like manner far from trivial [Ha and Schmidhuber(2018)]. Encoding high-dimensional observations (e.g., images) through commonplace embedding methods (e.g., through autoencoders or encoder-only transformers) leaves the latent representation generally uninterpretable and task-agnostic. This raises concerns about whether spurious correlations distort the latent space or if it effectively encodes the details necessary for discriminating between features that should remain functionally disentangled.
@@
Invariance and equivariance relations can help address uninterpretability in representation learning. These terms characterize how representations respond to observation-space transformations. If the representation of shifts in an expected manner due to a transformation , then the representation model is said to be equivariant to that transformation. Likewise, if the representation does not shift under the transformation, then the model is said to be invariant to the transformation. For example, bisimulation metrics help learn latent obstacle representations invariant to changes in type, size, and brightness [Zhang et al.(2020)Zhang, Lyle, Sodhani, Filos, Kwiatkowska, Pineau, Gal, and Precup]. [Pol et al.(2020)Pol, Kipf, Oliehoek, and Welling] use contrastive loss to enforce action equivariance. Yet, integrating expert priors remains difficult; one method maps complex observation transformations to simpler latent ones via a symmetric embedding network [Park et al.(2022)Park, Biza, Zhao, Meent, and Walters].
@@
We categorize representations along two dimensions: (1) the nature of their transformation response (invariance versus equivariance) and (2) their degree of human alignment (aligned versus misaligned). A representation that is aligned-invariant remains unchanged when an observation undergoes a meaning-preserving transformation, while an aligned-equivariant representation transforms predictably when the observation’s meaning is altered. In contrast, a representation is misaligned-invariant if it does not change under meaningful effects made to the observation (suggesting underfitting), and it is misaligned-equivariant if it changes in response to an observation transformation that should not affect the underlying meaning (suggesting a domain shift). Our training objective is to achieve invariance and equivariance alignment by ensuring that the post-transformation representations accurately reflect our human interpretation of the change.
Definition 3.3 (Equivariance).
@@
Let be an observation space transformation randomly parameterized by , and let be a latent space transformation randomly parameterized by . An encoder is an equivariant function if . Invariance is a special case of equivariance where returns its argument.
@@
where is the overall WM training loss, which is a function of the input observation .
Example. Consider a vision-based autonomous car that hands over its neural-based controls to a simpler safety controller if a collision is predicted by its world model, consisting of a “physical” and “style” branch per Principle 1. Based on the prior knowledge, an engineer decides that scene brightness should not affect the physical latents; hence, the physics encodings should be invariant to changes in observation brightness. However, the style encodings should represent brightness in the resulting latent embedding and, thus, should be equivariant to the changes in brightness.
3.3 Principle 3: Multi-Level and Multi-Strength Supervision for Latent Representations
To bridge rich observations and physical meaning, world models must adapt to supervision signals of varying form and strength [Lee et al.(2013), Chen et al.(2020)Chen, Kornblith, Norouzi, and Hinton] — from exact state labels to trajectory-level constraints and weak self-supervision. They must also integrate these signals based on abstraction level (e.g., exact values, intervals, or missing data) to align representations with physical systems. Multi-level supervision tailors the loss functions and training process to the level of abstraction (e.g., full trajectories vs. specific state dimensions) and strength (e.g., exact labels vs. intervals). For instance, physical state labels allow for the direct alignment of latent representations with real-world quantities using supervised loss. When such labels are unavailable, temporal consistency and smoothness of trajectories can serve as implicit regularization techniques to constrain learned representations. Finally, self-supervision can leverage data-driven structures to discover meaningful latent representations in entirely unsupervised settings.
Supervised Learning: Strong supervision directly aligns specific latent dimensions with known physical states (e.g., positions, velocities), enabling fine-grained interpretability. In many cases, supervision signals are introduced directly into the embeddings to capture key features from labeled data [Zhuang et al.(2015)Zhuang, Cheng, Luo, Pan, and He]. For example, in low-dimensional systems with position and velocity states , additional latent dimensions () can improve reconstruction quality and stability [Chen et al.(2016)Chen, Duan, Houthooft, Schulman, Sutskever, and Abbeel, Alemi et al.(2018)Alemi, Poole, Fischer, Dillon, Saurous, and Murphy, Rezende et al.(2014)Rezende, Mohamed, and Wierstra].